Assuming a velocity distribution which is a

parabola having its vertex 12 in from the boundary,

calculate the shear stress at y= 0, 3, 6, 9 and 12 inches.

Fluid’s absolute viscosity is 600 P


1
Expert's answer
2022-02-28T10:05:57-0500

"600~P=0.6~\frac{N\cdot s}{m^2}=0.01253~\frac{lb\cdot s}{ft^2},"

"120-u=a(12-y)^2,"

"u=0,~t=0,\implies a=\frac 56,"

"u=120-\frac 56(12-y)^2,"

"\frac{du}{dy}=\frac 53(12-y),"

"\def\arraystretch{1.5}
 \begin{array}{|c|c|c|c|c|c|}
\hline
 y & 0 & 3&6&9&12 \\ \hline
 \frac{du}{dy} & 20 & 15&10&5&0\\
 \hline
 \eta & 0.251&0.188&0.1253&0.0627 & 0
\\
\hline
\end{array}"